OpenStudy (anonymous):

Which property would you use to solve 6x = 42? Would it be multiplication property of equality?

OpenStudy (anonymous):

Division property

OpenStudy (anonymous):

no..it is division

OpenStudy (anonymous):

Division property isn't one of my choices. I can only choose between symmetric property of equality, transitive property of equality, multiplication property of equality, or addition property of equality.

OpenStudy (anonymous):

x=42/6=7.

OpenStudy (anonymous):

I know what the answer is, but I don't know which property of equality it belongs to.

OpenStudy (anonymous):

Division property.

OpenStudy (anonymous):

That's not one of my answers. I thought it was multiplication since you could add to both sides of the equation but I wasn't sure.

OpenStudy (anonymous):

What are your answer choices? List them here and I'll try to analyze them.

OpenStudy (anonymous):

symmetric property of equality, transitive property of equality, multiplication property of equality, or addition property of equality.

OpenStudy (anonymous):

You're right then. It should be the multiplication one because the other ones don't make sense at all.

Directrix (directrix):

Would it be multiplication property of equality? Yes, it could be that. 6x = 42 Multiply both sides by 1/6 (1/6)*(6x) = (1/6)*42) x = 7 Dividing by 6 is the same operation as multiplying by 1/6. @haileemackk

OpenStudy (anonymous):

Thank you! I just wanted to make sure. Thanks for explaining it.

Directrix (directrix):

You are welcome.